OpenStudy (anonymous):

Solve. Please show all of your work. You are traveling to your mother’s house that is 234 miles away. If you are currently twice as far from your home as you are from your mother’s house, how far have you traveled? Round your answer to the nearest tenth mile.

OpenStudy (anonymous):

Let distance from mother house =x then, distance traveled by me (i.e distance from my home)= 2x x + 2x = 234 [According to question] 3x= 234 => x = 234/3= 78 distance traveled by me = 2* 78 = 156 miles

OpenStudy (anonymous):

And rounded to the nearest tenth mile would be 156.0 miles